# A Survey on 3D Gaussian Splatting Applications: Segmentation, Editing, and Generation

### v3

@@ Abstract
In the context of novel view synthesis, 3D Gaussian Splatting (3DGS) has recently emerged as an efficient and competitive counterpart to Neural Radiance Field (NeRF), enabling high-fidelity photorealistic rendering in real time. Beyond novel view synthesis, the explicit and compact nature of 3DGS enables a wide range of downstream applications that require geometric and semantic understanding. This survey provides a comprehensive overview of recent progress in 3DGS applications. It first introduces 2D foundation models that support semantic understanding and control in 3DGS applications, followed by a review of NeRF-based methods that inform their 3DGS counterparts. We then categorize 3DGS applications into three foundational tasks: segmentation, editing, and generation, alongside additional functional applications built upon or tightly coupled with these foundational capabilities. For each, we summarize representative methods, supervision strategies, and learning paradigms, highlighting shared design principles and emerging trends. Commonly used datasets and evaluation protocols are also summarized, along with comparative analyses of recent methods across public benchmarks. To support ongoing research and development, a continually updated repository of papers, code, and resources is maintained at https://github.com/heshuting555/Awesome-3DGS-Applications.
@@ 1 Introduction
§1§2§3§3.1§3.2§3.3§4§5§6§A§2.1§2.2§2.3§3.1.1§3.1.2§3.1.3§3.1.4§3.2.1§3.2.2§3.2.3§3.3.1§3.3.2§A.1§A.2
@@
$\bullet$ Contribution. This survey provides a comprehensive and structured review of recent literature on 3DGS with a focus on its emerging downstream applications. We begin by introducing 2D foundation models that enable semantic control and understanding in 3DGS-based systems, and briefly revisit relevant NeRF-based methods to establish conceptual continuity. The core of our survey is organized around three foundational task categories: segmentation, editing, and generation. Within each category, we systematically compare representative works in terms of technical designs, learning paradigms, and supervision strategies. We further analyze benchmark settings and performance results in each domain, aiming to provide a self-contained and accessible resource for both newcomers and experienced researchers. Finally, we discuss open challenges and future research opportunities, with the goal of fostering deeper exploration and broader adoption of 3DGS across high-level vision and graphics application tasks.
Fig. 2 provides an overview of the structure of this survey. In Sec. 2, we introduce essential background knowledge, including key definitions, commonly used 2D foundation models, and relevant NeRF-based research that informs 3DGS development. Sec. 3 presents a task-centric review of recent 3DGS-based methods, with a focus on segmentation, editing, and generation. In Sec. 4, we summarize evaluation protocols and benchmark settings, and provide quantitative comparisons across tasks and datasets to assess the performance of existing approaches. Sec. 5 discusses open challenges and outlines potential future research directions. Sec. 6 concludes the survey with a summary of key insights and takeaways. Finally, we present a discussion of functional applications and complete dataset statistics tables in the Appendix.
@@ 2 Background
In this section, we first categorize the main research directions of 3DGS for downstream applications. We then provide a concise overview of widely used 2D foundation models that support these tasks. Finally, we review related research domains in the NeRF literature that lay the groundwork for 3DGS development.
@@ 2.1.1 Application Tasks Category
$\bullet$ 3DGS Segmentation. Based on the input space $\mathcal{X}$, existing 3DGS segmentation methods can be broadly classified into two categories: generic segmentation and promptable segmentation. Generic segmentation aims to predict pixel- or point-level labels without external input prompts, and can be further divided into: (i) Semantic segmentation [17]; (ii) Instance segmentation [18, 19, 20]; and (iii) Panoptic segmentation [21, 22]. Promptable segmentation extends the generic setting by introducing external input prompts to guide the segmentation process. Depending on the form of prompts, this category includes: (i) Interactive segmentation, where users provide spatial hints such as clicks, scribbles, boxes, or masks [23, 24, 25]; (ii) Open-vocabulary segmentation, which uses category names as textual prompts to support segmenting novel or unseen classes [26, 27]; (iii) Referring segmentation, which relies on natural language expressions to localize and segment specific objects [7]. In addition, depending on the representation of the output space $\mathcal{Y}$, segmentation results can be produced either on rendered 2D images [26, 28] or directly over the 3D Gaussian representations [29, 30, 31, 32].
$\bullet$ 3DGS Editing. Editing tasks aim to modify scene attributes such as geometry, appearance, or illumination, while maintaining structural and visual consistency across viewpoints. In this setting, $\mathcal{Y}$ denotes the edited scene state, often represented by an updated set of Gaussians. 3DGS facilitates localized and differentiable editing operations, including addition, deletion, and transformation of scene components, thus enabling applications such as content-aware inpainting and style-consistent relighting. Based on the form of inputs $\mathcal{X}$, editing approaches can be categorized into image-guided [33, 34] and text-guided methods [8, 35, 36].
@@ 3.1.1 Feature Distillation-based Methods
$\bullet$ Enhancing Segmentation Accuracy. N2F2 [88] addresses LangSplat’s inefficiency in scale selection by learning a unified high-dimensional feature field under hierarchical supervision, where each feature dimension encodes scene semantics at different granularities, enabling fine-grained yet efficient representation without multi-scale querying. LangSurf [89] derives dense pixel-level semantics with SAM pooling, and jointly optimizes surface language Gaussians with geometric supervision and contrastive loss for improved 3D spatial consistency. seconGS [90] refines LangSplat with SAM2 masklet supervision and a two-step querying scheme that first retrieves distilled ground-truth and then queries individual Gaussians for more accurate 3D understanding. VLGaussian [28] observes that prior color-based rasterization is unsuitable for language modalities due to its dependence on the same color opacity. Instead, it introduces a cross-modal rasterizer tailored for rendering semantic language features with newly introduced opacity for language. SuperGSeg [91] clusters Gaussians into Super-Gaussians with spatial and instance cues built on Scaffold-GS [3] and assigns language features per cluster for structured scene understanding. FreeGS [92] introduces a semantic-embedded 3DGS framework by coupling semantic features with instance identities via the proposed IDSF field. A two-step alternating optimization and a 2D–3D contrastive loss ensure view-consistent semantics without requiring 2D labels. econSG [93] improves zero-shot 3D semantic understanding by projecting refined multi-view language features into a low-dimensional 3D latent space, enabling efficient optimization and consistent semantic field initialization within the 3D Gaussian framework. CCL-LGS [94] aligns SAM-generated masks across views via a zero-shot tracker, extracts semantics with CLIP, and applies contrastive codebook learning to enforce cross-view consistency by promoting intra-class compactness and inter-class separation. ILGS [95] mitigates view-inconsistent language embeddings by enforcing cross-view semantic alignment with an identity-aware semantic consistency loss and refining boundaries via progressive mask expansion. COS3D [96] learns collaborative instance and language fields, using SAM masks with contrastive instance supervision and an instance-to-language mapping. At inference, it uses adaptive language-to-instance prompt refinement to produce accurate output. Beyond the aforementioned approaches that solely introduce semantic information, GLS [97] incorporates geometric priors (surface normals) to effectively refine the rendered depth.
@@
$\bullet$ Training-Free Mask Lifting. To avoid the computational cost and data requirements of model training, several methods explore training-free strategies that directly associate 2D segmentation masks with 3D Gaussian primitives for efficient 3D segmentation. SAGD [23] adopts a simple yet effective projection-based strategy: the centers of 3D Gaussians are projected onto 2D masks, and those falling within foreground regions are classified as foreground Gaussians. FlashSplat [119] formulates the mask lifting process as a one-step linear programming (LP) optimization problem, directly transforming 2D mask supervision into 3D Gaussian masks without iterative training. GaussianCut [29] represents the scene as a graph over Gaussians and applies graph-cut optimization to partition them into foreground and background. It leverages coarse segmentation from 2D image/video models and refines the results through learned edge affinities in the constructed graph. THGS [120] constructs a hierarchical superpoint graph from 3D Gaussian primitives and reprojects 2D semantic features onto this structure, producing a view-consistent semantic field without the need for supervised training. iSegMan[25] introduces a visibility-guided voting scheme that links 2D segmentations from SAM with 3D Gaussians, treating the association as a voting process weighted by Gaussian visibility. LBG [121] assigns semantics to 3D Gaussians via a 2D-to-3D lifting and incrementally merges them based on semantic and geometric overlap, enabling fast training-free 3D instance segmentation. LUDVIG [122] introduces “inverse rendering” aggregation, where 2D features or masks from multiple views are uplifted to the 3D by performing a weighted average guided by the 3DGS rendering weights. Then, a graph diffusion mechanism is designed for feature refinement.
@@ 3.2.3 Style Transfer
$\bullet$ Efforts toward Integrating Diffusion Model. InstantStyleGaussian [175] accelerates style editing by using a diffusion model to generate target style images, incorporating them into the training set, and iteratively optimizing the Gaussian splatting scene. ArtNVG [176] enhances control over content and style via the CSGO model and Tile ControlNet, and introduces an Attention-based Neighboring-View Alignment mechanism to maintain consistent colors and textures across neighboring views. Morpheus [177] employs an RGBD diffusion model with depth-guided cross-attention, feature injection, and a Warp ControlNet conditioned on composite frames to guide 3D stylization. FantasyStyle [178] proposes Controllable Stylized Distillation, achieving 3D style transfer solely through diffusion model distillation.
@@
$\bullet$ Feed-Forward-based Model. Styl3R [183] introduces a branched architecture that separates structural modeling from appearance shading, effectively preventing style transfer from distorting the underlying 3D scene structure, while employing an identity loss to facilitate training of the stylized model. $\text{A}^{3}$GS [184], a novel feed-forward framework, introduces an AdaIN-based stylizer that effectively injects the style features of a target image into the latent representations of a 3DGS scene.

### v4

@@ Abstract
In the context of novel view synthesis, 3D Gaussian Splatting (3DGS) has recently emerged as an efficient and competitive counterpart to Neural Radiance Field (NeRF), enabling high-fidelity photorealistic rendering in real time.Beyond novel view synthesis, the explicit and compact nature of 3DGS enables a wide range of downstream applications that require geometric and semantic understanding. This survey provides a comprehensive overview of recent progress in 3DGS applications. It first reviews the reconstruction preliminaries of 3DGS, followed by the problem formulation, 2D foundation models, and related NeRF-based research areas that inform downstream 3DGS applications. We then categorize 3DGS applications into three foundational tasks: segmentation, editing, and generation, alongside additional functional applications built upon or tightly coupled with these foundational capabilities. For each, we summarize representative methods, supervision strategies, and learning paradigms, highlighting shared design principles and emerging trends. Commonly used datasets and evaluation protocols are also summarized, along with a comparative analysis of recent methods across public benchmarks. To support ongoing research and development, a continually updated repository of papers, code, and resources is maintained at https://github.com/heshuting555/Awesome-3DGS-Applications.
@@ 1 Introduction
Although presented separately for clarity, these three tasks are closely interconnected. Segmentation provides semantic grounding and object-level decomposition, which are often prerequisites for precise and localized editing, such as modifying or removing a target object. Generation further extends this pipeline from modifying existing content to synthesizing new objects or scenes. Conversely, generative priors and synthetic data can in turn improve segmentation and editing quality, forming a mutually reinforcing loop. Rather than three parallel directions, these three tasks are deeply intertwined.
@@
$\bullet$ Contribution. This survey provides a comprehensive and structured review of recent literature on 3DGS with a focus on its emerging downstream applications. We first review the reconstruction preliminaries of 3DGS, together with relevant 2D foundation models and related NeRF-based research areas, to establish conceptual continuity. The core of our survey is organized around three foundational task categories: segmentation, editing, and generation. Within each category, we systematically compare representative works in terms of technical designs, learning paradigms, and supervision strategies. We further analyze benchmark settings and performance results in each domain, aiming to provide a self-contained and accessible resource for both newcomers and experienced researchers. Finally, we discuss open challenges and future research opportunities, with the goal of fostering deeper exploration and broader adoption of 3DGS across high-level vision and graphics application tasks.
Fig. 2 provides an overview of the structure of this survey. In Sec. 2, we review the essential background, including the reconstruction preliminaries of 3DGS, the problem formulation and taxonomy of downstream applications, commonly used 2D foundation models, and related NeRF-based research areas that inform 3DGS development. Sec. 3 presents a task-centric review of recent 3DGS-based methods, with a focus on segmentation, editing, and generation. Sec. 4 discusses open challenges and outlines potential future research directions. Sec. 5 concludes the survey with a summary of key insights and takeaways. Finally, we provide additional discussions of performance comparisons among recent methods and functional application tasks in the Appendix.
@@ 2 Background
In this section, we first review the preliminaries of 3DGS-based reconstruction, and then outline the main research directions for downstream applications. We further provide a concise overview of widely used 2D foundation models that support these tasks, and finally discuss related research domains in the NeRF literature that have laid the foundation for the development of 3DGS.
2.1 Preliminaries in 3DGS: Reconstruction
Although this survey focuses on downstream application tasks such as segmentation, editing, and generation, reconstruction and novel view synthesis (NVS) remain the historical and technical foundation of the broader 3DGS ecosystem. Revisiting these reconstruction- and representation-oriented foundations helps clarify the technical basis upon which downstream applications are built. Below, we briefly summarize several major upstream directions that provide important context for, yet remain complementary to, the main focus of this survey.
$\bullet$ Vanilla 3DGS Reconstruction. The original 3D Gaussian Splatting (3DGS) framework [1] is introduced for high-fidelity 3D reconstruction and real-time novel-view synthesis. It represents a scene as a set of explicit 3D Gaussian primitives $\mathcal{G}=\{g_{i}\}_{i=1}^{\mathcal{N}}$, where $\mathcal{N}$ is the number of Gaussians. Each Gaussian $g_{i}$ is parameterized by its mean position $\mu_{i}\in\mathbb{R}^{3}$, covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$, opacity $\sigma_{i}\in\mathbb{R}^{1}$, and color $c_{i}\in\mathbb{R}^{d_{c}}$, where $d_{c}=3$ for RGB color parameters. To render an image, each 3D Gaussian is projected onto the 2D image plane, where it contributes to the pixel color based on its opacity. The color at a given pixel $v$, denoted as $C(v)$, is computed by blending the contributions of all Gaussians according to their opacity. This process is formulated as:
$C(v)=\sum_{i=1}^{\mathcal{N}}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (1)
where $c_{i}$ is the color vector of the $i$-th Gaussian, and $\alpha_{i}=\sigma_{i}G^{2D}_{i}(v)$, $G^{2D}_{i}(v)$ is the 2D projection function. This explicit representation renders substantially faster than NeRF-style implicit pipelines while preserving strong reconstruction quality. However, vanilla 3DGS is mainly tailored to dense-view, static, and well-posed scenes, motivating many subsequent extensions.
$\bullet$ Sparse-view Reconstruction. An important line of research studies how to extend 3DGS from dense-view settings to sparse-view or few-shot reconstruction, where limited observations often lead to geometric degradation, overfitting to seen views, and unstable novel-view synthesis [17, 18, 19]. Existing methods mainly follow two directions. The first is regularization-based reconstruction, which introduces additional geometric cues such as monocular or multi-view depth priors, smoothness constraints, cost-volume guidance, or Gaussian densification strategies to improve surface completeness and global consistency under sparse inputs [20, 18, 17, 19, 4, 21]. The second direction leverages stronger learned priors, either by directly inferring Gaussian representations from sparse observations or by augmenting sparse inputs with generative models [22, 23, 24, 25, 26].
$\bullet$ Generalizable 3DGS. To overcome the limitation of per-scene optimization in vanilla 3DGS, a growing body of work studies generalizable 3DGS, which aims to predict Gaussian representations for unseen scenes directly from one or multiple input images using learned priors. Existing methods can be broadly grouped by the form of the predicted representation. One line predicts Gaussian maps, where each image location encodes a 3D Gaussian, typically using encoder-decoder or Transformer-based architectures for pixel-aligned Gaussian prediction [27, 24, 28, 29, 30, 31]. To improve geometric quality, many methods further incorporate multi-view geometric reasoning, such as epipolar constraints and cost-volume representations, for more reliable depth and cross-view correspondence estimation [22, 32, 33, 4, 34, 35]. Another line predicts more structured Gaussian volumes or triplane-based representations, which provide stronger 3D inductive bias for feed-forward reconstruction [36, 37, 38, 39, 40, 41]. In addition, several methods leverage pre-trained reconstruction or generative foundation models as geometry or appearance priors, further improving generalization under sparse-view or pose-free settings [42, 43, 44, 45, 46].
$\bullet$ Efficiency and Compression. Although 3DGS enables high-quality real-time rendering, its practical deployment still faces considerable memory and storage challenges due to the large number of Gaussians and their associated attributes. Existing efficiency-oriented studies mainly improve compactness from two perspectives. The first is to reduce the number of Gaussians through pruning, importance-based selection, or more structured representations such as anchor-, grid-, or octree-based designs [3, 47, 48, 49, 50, 51, 52]. The second is to compress Gaussian attributes by using vector quantization, codebooks, residual quantization, or other compact storage schemes for color, spherical harmonics, scale, rotation, and related properties [53, 47, 54, 48, 55]. Some methods further combine both strategies by coupling structured representations with adaptive quantization to reduce redundancy while maintaining rendering quality [49, 3].
$\bullet$ Dynamic Reconstruction. Beyond static scene reconstruction, 3DGS has also been extended to dynamic 3D reconstruction, where the core challenge is to model time-varying scene content while preserving geometric fidelity and temporal consistency. In general, dynamic reconstruction methods typically need to address two key issues: how to separate static and dynamic components in the scene, and how to represent the motion of dynamic objects. Existing methods mainly follow three paradigms. The first is time-varying modeling, which directly parameterizes Gaussian attributes such as position, rotation, opacity, or appearance as functions of time [56, 57, 58, 59, 60, 61]. The second is deformation-based modeling, which defines a canonical static space and learns a deformation field to warp Gaussians across time [62, 63, 64, 65, 66, 67]. The third is 4D Gaussian modeling, which treats spacetime as a unified entity by extending 3D Gaussians into 4D primitives [68, 69]. These methods have enabled dynamic reconstruction in a wide range of scenarios, from object- and human-centric settings to large-scale driving and urban environments [70, 60, 57].
$\bullet$ Mesh Extraction and Surface Recovery. Beyond view synthesis, reconstruction-oriented 3DGS research has also explored mesh extraction and explicit surface recovery, which are important for applications requiring explicit geometry, physical interaction, or geometry-aware editing. Existing methods typically combine Gaussian representations with surface regularization or hybrid geometric formulations to improve the recoverability of explicit surfaces [71, 72]. In this setting, evaluation involves not only rendering fidelity but also geometric quality criteria such as surface accuracy, completeness, smoothness, and topology preservation.
$\bullet$ Photorealism and Quality Enhancement. Beyond efficient novel view synthesis, another important research direction aims to improve the visual fidelity of 3DGS rendering. Existing studies mainly enhance photorealism from three aspects. The first is to reduce aliasing and splatting artifacts, which often arise from the discrete sampling and projection process of Gaussian splatting, especially under varying resolutions or focal lengths [73, 74, 75, 76]. The second is to improve view-dependent appearance, including reflections, specularities, and relightable effects, by incorporating more expressive lighting and shading models, inverse rendering formulations, or physically inspired appearance representations [77, 78, 79, 80, 81, 82]. The third is to address real-world degradations, such as motion blur, defocus blur, and unstable visibility, which can significantly reduce rendering realism in practical captures [83, 84, 85].
$\bullet$ Physically Grounded 3D Modeling. Beyond geometrically plausible reconstruction and rendering, an emerging direction aims to make 3DGS physically grounded, so that dynamic scene evolution is not only visually convincing but also consistent with underlying physical behaviors. Existing works typically build on dynamic or 4D Gaussian representations [63, 68, 66, 86, 87, 65] and further introduce physics-aware mechanisms in two main ways. One line incorporates explicit simulation priors, such as material point methods, position-based dynamics, or other mechanics-inspired simulators, by treating Gaussians as particles or material carriers for deformation and motion [88, 89, 90, 91, 92, 93, 94, 95, 96]. Another line imposes physically motivated constraints on motion and deformation, such as rigidity, isometry, momentum conservation, to encourage more faithful and temporally consistent dynamics [86, 87].
@@ 2.2.1 Application Tasks Category
$\bullet$ 3DGS Segmentation. Based on the input space $\mathcal{X}$, existing 3DGS segmentation methods can be broadly classified into two categories: generic segmentation and promptable segmentation. Generic segmentation aims to predict pixel- or point-level labels without external input prompts, and can be further divided into: (i) Semantic segmentation [97]; (ii) Instance segmentation [98, 99, 100]; and (iii) Panoptic segmentation [101, 102]. Promptable segmentation extends the generic setting by introducing external input prompts to guide the segmentation process. Depending on the form of prompts, this category includes: (i) Interactive segmentation, where users provide spatial hints such as clicks, scribbles, boxes, or masks [103, 104, 105]; (ii) Open-vocabulary segmentation, which uses category names as textual prompts to support segmenting novel or unseen classes [106, 107]; (iii) Referring segmentation, which relies on natural language expressions to localize and segment specific objects [7]. In addition, depending on the representation of the output space $\mathcal{Y}$, segmentation results can be produced either on rendered 2D images [106, 108] or directly over the 3D Gaussian representations [109, 110, 111].
$\bullet$ 3DGS Editing. Editing tasks aim to modify scene attributes such as geometry, appearance, or illumination, while maintaining structural and visual consistency across viewpoints. In this setting, $\mathcal{Y}$ denotes the edited scene state, often represented by an updated set of Gaussians. 3DGS facilitates localized and differentiable editing operations, including addition, deletion, and transformation of scene components, thus enabling applications such as content-aware inpainting and style-consistent relighting. Based on the form of inputs $\mathcal{X}$, editing approaches can be categorized into image-guided [112, 113] and text-guided methods [8, 114].
@@ 3.1.1 Feature Distillation-based Methods
$\bullet$ Enhancing Segmentation Accuracy. N2F2 [165] addresses LangSplat’s inefficiency in scale selection by learning a unified high-dimensional feature field under hierarchical supervision, where each feature dimension encodes scene semantics at different granularities, enabling fine-grained yet efficient representation without multi-scale querying. LangSurf [166] derives dense pixel-level semantics with SAM pooling, and jointly optimizes surface language Gaussians with geometric supervision and contrastive loss for improved 3D spatial consistency. seconGS [167] refines LangSplat with SAM2 masklet supervision and a two-step querying scheme that first retrieves distilled ground-truth and then queries individual Gaussians for more accurate 3D understanding. VLGaussian [108] observes that prior color-based rasterization is unsuitable for language modalities due to its dependence on the same color opacity. Instead, it introduces a cross-modal rasterizer tailored for rendering semantic language features with newly introduced opacity for language. SuperGSeg [168] clusters Gaussians into Super-Gaussians with spatial and instance cues built on Scaffold-GS [3] and assigns language features per cluster for structured scene understanding. FreeGS [169] introduces a semantic-embedded 3DGS framework by coupling semantic features with instance identities via the proposed IDSF field. A two-step alternating optimization and a 2D–3D contrastive loss ensure view-consistent semantics without requiring 2D labels. econSG [170] improves zero-shot 3D semantic understanding by projecting refined multi-view language features into a low-dimensional 3D latent space, enabling efficient optimization and consistent semantic field initialization within the 3D Gaussian framework. CCL-LGS [171] aligns SAM-generated masks across views via a zero-shot tracker, extracts semantics with CLIP, and applies contrastive codebook learning to enforce cross-view consistency by promoting intra-class compactness and inter-class separation. ILGS [172] mitigates view-inconsistent language embeddings by enforcing cross-view semantic alignment with an identity-aware semantic consistency loss and refining boundaries via progressive mask expansion. COS3D [173] learns collaborative instance and language fields, using SAM masks with contrastive instance supervision and an instance-to-language mapping. At inference, it uses adaptive language-to-instance prompt refinement to produce accurate output. Beyond the aforementioned approaches that solely introduce semantic information, GLS [174] incorporates geometric priors (surface normals) to effectively refine the rendered depth. REALM [175] enables reasoning-based open-world 3D segmentation via a global-to-local MLLM grounding strategy, avoiding extensive 3D-specific post-training while supporting downstream 3D interaction tasks.
@@
$\bullet$ Training-Free Mask Lifting. To avoid the computational cost and data requirements of model training, several methods explore training-free strategies that directly associate 2D segmentation masks with 3D Gaussian primitives for efficient 3D segmentation. SAGD [103] adopts a simple yet effective projection-based strategy: the centers of 3D Gaussians are projected onto 2D masks, and those falling within foreground regions are classified as foreground Gaussians. FlashSplat [198] formulates the mask lifting process as a one-step linear programming (LP) optimization problem, directly transforming 2D mask supervision into 3D Gaussian masks without iterative training. GaussianCut [109] represents the scene as a graph over Gaussians and applies graph-cut optimization to partition them into foreground and background. It leverages coarse segmentation from 2D image/video models and refines the results through learned edge affinities in the constructed graph. THGS [199] constructs a hierarchical superpoint graph from 3D Gaussian primitives and reprojects 2D semantic features onto this structure, producing a view-consistent semantic field without the need for supervised training. ReLaGS [200] builds upon THGS by constructing an open-vocabulary 3D semantic scene graph for efficient 3D perception and relational reasoning without scene-specific training. iSegMan[105] introduces a visibility-guided voting scheme that links 2D segmentations from SAM with 3D Gaussians, treating the association as a voting process weighted by Gaussian visibility. LBG [201] assigns semantics to 3D Gaussians via a 2D-to-3D lifting and incrementally merges them based on semantic and geometric overlap, enabling fast training-free 3D instance segmentation. LUDVIG [202] introduces “inverse rendering” aggregation, where 2D features or masks from multiple views are uplifted to the 3D by performing a weighted average guided by the 3DGS rendering weights. Then, a graph diffusion mechanism is designed for feature refinement.
@@ 3.2.3 Style Transfer
$\bullet$ Efforts toward Integrating Diffusion Model. InstantStyleGaussian [256] accelerates style editing by using a diffusion model to generate target style images, incorporating them into the training set, and iteratively optimizing the Gaussian splatting scene. ArtNVG [257] enhances control over content and style via the CSGO model and Tile ControlNet, and introduces an Attention-based Neighboring-View Alignment mechanism to maintain consistent colors and textures across neighboring views. Morpheus [258] employs an RGBD diffusion model with depth-guided cross-attention, feature injection, and a Warp ControlNet conditioned on composite frames to guide 3D stylization. FantasyStyle [259] proposes Controllable Stylized Distillation, achieving 3D style transfer solely through diffusion model distillation. DiffStyle3D [260] performs style transfer by directly computing the distance to the style image within the self-attention feature space of the diffusion model.
@@
$\bullet$ Feed-Forward-based Model. Styl3R [265] introduces a branched architecture that separates structural modeling from appearance shading, effectively preventing style transfer from distorting the underlying 3D scene structure, while employing an identity loss to facilitate training of the stylized model. $\text{A}^{3}$GS [266], a novel feed-forward framework, introduces an AdaIN-based stylizer that effectively injects the style features of a target image into the latent representations of a 3DGS scene. Stylos [267] uses a Transformer with two paths: self-attention for accurate geometry, and cross-attention to apply consistent style across views.